# Mask-Adapter: The Devil is in the Masks for Open-Vocabulary Segmentation

### v1

@@ 1 Introduction
Image segmentation, which aims to assign each pixel a label from a predefined set of categories, has seen significant advancements. Despite the success of numerous image segmentation approaches [6, 5, 40, 36, 48, 33], these approaches cannot segment categories unseen during training, limiting their applicability in open-world scenarios. Open-vocabulary segmentation (OVS) has expanded traditional image segmentation from fixed-set to open-set, enabling the segmentation of categories not appeared in the training set. In contrast to fixed-category segmentation, open-vocabulary segmentation not only supports open-set segmentation but also enables challenging segmentation based on arbitrary text inputs, e.g., categories and descriptions.
@@
In this paper, we propose a simple yet effective method to address the bottlenecks in open-vocabulary segmentation. Instead of directly using predicted masks and mask pooling (or mask cropping) to extract mask embeddings, we introduce a Mask-Adapter (Fig. 2 (c)) to extract semantic activation maps from proposal masks and CLIP features. Mask embeddings are aggregated based on both semantic and contextual information and subsequently matched with text embeddings for mask recognition. Fig. 2 compares mask cropping, mask pooling, and Mask-Adapter methods. In contrast to previous approaches, Mask-Adapter offers several key advantages: (1) rather than neglecting the background, it aggregates mask embeddings across the entire image, incorporating contextual information to enrich the feature representation; (2) unlike mask pooling, which only conveys the positional information of target regions, semantic activation maps selectively highlights informative regions relevant to recognition, while suppressing less informative areas, thus enhancing feature discriminability; (3) our method preserves the generalization capabilities of CLIP during training while concurrently improving mask recognition performance. Furthermore, we introduce a mask consistency loss to enforce similar proposal masks to obtain similar CLIP mask embeddings, thereby improving robustness to variations in predicted masks. To further mitigate overfitting, we replace the Hungarian matcher with an IoU-based matcher, enabling the model to train on a broader range of proposal masks. In addition, we propose a mixed-mask training strategy that enhances the robustness of the Mask Adapter and improves its performance on open-vocabulary segmentation tasks.
@@ 2.1 Open-Vocabulary Segmentation
Open-vocabulary segmentation aims to segment images based on arbitrary textual descriptions rather than predefined classes, primarily categorizing into two types: per-pixel classification [20, 8, 39] and mask classification [14, 42, 9, 15, 22, 41, 46, 18]. OpenSeg [14] first employs mask pooling method to aggregate region features for mask classification. SimpleBaseline [42] and Zegformer [9] propose two-stage frameworks which generate class-agnostic masks and feed the masked image into CLIP for classification. Per-pixel methods, e.g., CAT-Seg [8] and SED [39], adopt cost aggregation to refine the pixel-text cost volume and predict per-pixel classification outputs. SAN [43] attaches an adapter to CLIP, enabling mask proposal generation and recognition. FC-CLIP [46] utilizes CLIP as a shared backbone for both mask generation and mask classification. However, both mask cropping and mask pooling approaches used in previous mask classification methods to obtain mask embeddings exhibit a limited upper bound performance, as illustrated in Fig. 1.
@@ 2.2 Adaptation for Vision-Language Models
To address CLIP’s limitations in dense prediction tasks, several methods have been proposed. CLIPSelf [37] and SILC [27] leverage self-distillation to align global and local features. Other approaches, such as MaskCLIP [50], CLIPSurgery [21], SCLIP [35], and GEM [1], modify CLIP’s attention structure to enhance its performance on dense prediction tasks. Furthermore, GBA [44] introduces learnable adapters to boost CLIP’s generalization in open-vocabulary segmentation, while PixelCLIP [31] improves CLIP’s fine-grained pixel representation by leveraging vision foundation models (e.g., DINO [4], SAM [19]) and unlabeled masks. Despite these advancements, CLIP continues to underperform on open-vocabulary region recognition tasks.
@@ 4.2 Main Results
Tab. 3 presents results of various mask embedding extraction methods on the ADE20K dataset using ground-truth masks. Our method achieves 74.1% accuracy, outperforming the methods in [42, 9, 22, 18, 46] demonstrating the effectiveness of our Mask-Adapter in transferring CLIP’s open-vocabulary recognition capabilities to mask classification.
4.3 Ablation Experiments
Ablation study on Mask-Adapter.
@@ B Additional Results
num. of maps GT mIoU FC-CLIP mIoU Adapter mIoU
1 41.9 34.6 32.9
16 43.4 35.4 34.1
Table 10: Comparison with different numbers of semantic activation maps during ground-truth warmup. GT mIoU and FC-CLIP mIoU represent results using Ground-truth and FC-CLIP predicted masks, respectively, while Adapter mIoU shows results with FC-CLIP predicted masks excluding the ensemble operation.

### v2

@@ 1 Introduction
Image segmentation, which aims to assign each pixel a label from a predefined set of categories, has seen significant advancements. Despite the success of numerous image segmentation approaches [6, 5, 43, 39, 52, 36, 7], these approaches cannot segment categories unseen during training, limiting their applicability in open-world scenarios. Open-vocabulary segmentation (OVS) has expanded traditional image segmentation from fixed-set to open-set, enabling the segmentation of categories not appeared in the training set. In contrast to fixed-category segmentation, open-vocabulary segmentation not only supports open-set segmentation but also enables challenging segmentation based on arbitrary text inputs, e.g., categories and descriptions.
@@
In this paper, we propose a simple yet effective method to address the bottlenecks in open-vocabulary segmentation. Instead of directly using predicted masks and mask pooling (or mask cropping) to extract mask embeddings, we introduce a Mask-Adapter (Fig. 2 (c)) to extract semantic activation maps from proposal masks and CLIP features, inspired by SparseInst [7]. Mask embeddings are aggregated based on both semantic and contextual information and subsequently matched with text embeddings for mask recognition. Fig. 2 compares mask cropping, mask pooling, and Mask-Adapter methods. In contrast to previous approaches, Mask-Adapter offers several key advantages: (1) rather than neglecting the background, it aggregates mask embeddings across the entire image, incorporating contextual information to enrich the feature representation; (2) unlike mask pooling, which only conveys the positional information of target regions, semantic activation maps selectively highlights informative regions relevant to recognition, while suppressing less informative areas, thus enhancing feature discriminability; (3) our method preserves the generalization capabilities of CLIP during training while concurrently improving mask recognition performance. Furthermore, we introduce a mask consistency loss to enforce similar proposal masks to obtain similar CLIP mask embeddings, thereby improving robustness to variations in predicted masks. To further mitigate overfitting, we replace the Hungarian matcher with an IoU-based matcher, enabling the model to train on a broader range of proposal masks. In addition, we propose a mixed-mask training strategy that enhances the robustness of the Mask Adapter and improves its performance on open-vocabulary segmentation tasks.
@@ 2.1 Open-Vocabulary Segmentation
Open-vocabulary segmentation aims to segment images based on arbitrary textual descriptions rather than predefined classes, primarily categorizing into two types: per-pixel classification [22, 10, 42] and mask classification [16, 45, 11, 17, 24, 44, 49, 20]. OpenSeg [16] first employs mask pooling method to aggregate region features for mask classification. SimpleBaseline [45] and Zegformer [11] propose two-stage frameworks which generate class-agnostic masks and feed the masked image into CLIP for classification. Per-pixel methods, e.g., CAT-Seg [10] and SED [42], adopt cost aggregation to refine the pixel-text cost volume and predict per-pixel classification outputs. PrimitiveNet [26] decomposes complex language constraints into simple primitives and employs cross-primitive and language-primitive attention mechanisms to enhance referring segmentation performance. SAN [46] attaches an adapter to CLIP, enabling mask proposal generation and recognition. FC-CLIP [49] utilizes CLIP as a shared backbone for both mask generation and mask classification. However, both mask cropping and mask pooling approaches used in previous mask classification methods to obtain mask embeddings exhibit a limited upper bound performance, as illustrated in Fig. 1.
@@ 2.2 Adaptation for Vision-Language Models
To address CLIP’s limitations in dense prediction tasks, several methods have been proposed. CLIPSelf [40] and SILC [30] leverage self-distillation to align global and local features. Furthermore, Yolo-World [8] integrates the CLIP text encoder with a lightweight detector for open-vocabulary object detection, while EVF-SAM [51] utilizes early vision-language fusion to enable text-prompted SAM for referring expression segmentation tasks. Other approaches, such as MaskCLIP [54], CLIPSurgery [23], SCLIP [38], and GEM [1], modify CLIP’s attention structure to enhance its performance on dense prediction tasks. Furthermore, GBA [47] introduces learnable adapters to boost CLIP’s generalization in open-vocabulary segmentation, while PixelCLIP [34] improves CLIP’s fine-grained pixel representation by leveraging vision foundation models (e.g., DINO [4], SAM [21]) and unlabeled masks. Despite these advancements, CLIP continues to underperform on open-vocabulary region recognition tasks.
@@ 4.2 Main Results
Tab. 3 presents results of various mask embedding extraction methods on the ADE20K dataset using ground-truth masks. Our method achieves 74.1% accuracy, outperforming the methods in [45, 11, 24, 20, 49] demonstrating the effectiveness of our Mask-Adapter in transferring CLIP’s open-vocabulary recognition capabilities to mask classification. Additionally, rows 2, 6, and 7 directly compare mask cropping, mask pooling, and Mask-Adapter using the same model (ConvNeXt-L) and source (OpenCLIP), with Mask-Adapter outperforming the others by 66.3% and 20.5%, respectively.
@@
4.3 Ablation Experiments
Ablation study on Mask-Adapter.
@@ B Additional Results
num. of maps GT mIoU FC-CLIP mIoU
1 41.9 34.6
16 43.4 35.4
Table 10: Comparison with different numbers of semantic activation maps during ground-truth warmup. GT mIoU and FC-CLIP mIoU represent results using Ground-truth and FC-CLIP predicted masks, respectively.
Ablations on the block structures.
Block GT mIoU FC-CLIP mIoU
ResNet Block 43.0 35.1
ConvNeXt Block 43.4 35.4
Transformer Block 43.4 35.0
Swin Transformer Block 42.6 34.9
Table 11: Ablation on different blocks structures.
In our paper, we primarily adopt ConvNeXt blocks, which are consistent with the backbone. We also provide the ablations about different blocks in Tab. 11. The results demonstrate that the ConvNeXt structure yields superior performance compared to other block configurations, and CNN-based blocks are generally more suitable for dense prediction tasks.